ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 1, 92 Behrends, Erhardt, FB Mathematik, Freie Universität Berlin; Nietzsch, Josef, Humboldt-Universität zu Berlin; Stroth, Gernot, Martin-Luther Universität Halle-Wittenberg The Study of Mathematics by High School Students at the University The project was carried out at two universities of Berlin from 1991 to 1997. Students of grade eleven and twelve took courses in mathematics over four terms that covered part of the mathematics curriculum for university students of mathematics. They attend weekly two-hour lectures and a severalhour tutorium. Besides that, they took part in an intensive course of one week and in several free lectures. In the fourth term, the lecture was substituted by a seminar in which the students had to study a mathematical problem in depth on their own and give a presentation of their work. At the end, they took a final exam.